Ammonium nitrate is widely used as a fertilizer, but it can also be used as an explosive. The decomposition of the solid ammonium nitrate produced nitrogen gas, oxygen gas and water vapour. Predict the numbers of molecules of water vapor that are produced along with 6.02 x 1023 molecules of nitrogen.
1
Expert's answer
2020-06-17T05:56:02-0400

Solution:

The balanced chemical equation:

2NH4NO3 → 2N2 + O2 + 4H2O

According to the chemical equation: n(N2)/2 = n(H2O)/4

Or: n(N2) = n(H2O)/2


One mole of any substance contains 6.02×1023 units of that substance.

Thus:

N(N2) = N(H2O)/2

N(H2O)= 2 × N(N2) = 2 × (6.02×1023) = 1.204×1024 molecules


Answer: 1.204×1024 molecules of water.